OpenStudy (anonymous):

The fractions -42/-45 and 14/15 are equivalent fractions.

hero (hero):

\[\frac{-42}{-45} = \frac{14}{15}\] \[\frac{42}{45} = \frac{14}{15}\] 42*15 = 14*45 3*14*15 = 3*14*15 True

OpenStudy (anonymous):

Alternate Solution: \[ \Large {\dfrac {-42}{-45} = \dfrac {42}{45} = \dfrac {3 \cdot 14}{3 \cdot 15} = \dfrac {14}{15}}. \]